A simple harmonic oscillator has an amplitude of 3.50 cm and a maximum speed of 26.0 cm/s. What is its speed when the displacement is 1.75 cm?
1 : 12.0 cm/s
2 : 22.5 cm/s
3 : 14.2 cm/s
4 : 15.0 cm/s
5 : 17.0 cm/s
1
Expert's answer
2020-08-02T15:07:57-0400

"X_m=3.50"


"\omega X_m=26cm/s"


"\omega=\frac{26}{3.5}=7.43"


"v=\omega \sqrt {X^2_m-X}"


"8\sqrt{0.035^2-0.017^2} =0.225"


which is "22.5cm\/s"